OpenStudy (anonymous):

Solve using law of cosines a = 12, b = 14, c = 21

OpenStudy (anonymous):

Do you know the formula for law of cosines?

OpenStudy (anonymous):

yea i did it but i wanna know if i got it right 12^2 = 14^2 + 21^2 – 2(14)(21) cosA 144 = 196 + 441 – 588 cosA 144 = 637 – 588 cosA -637 -637 -493 = -588 cosA -588 -493/-588 = cosA cos^-1(493/588) = 33.02

OpenStudy (anonymous):

That looks correct to me.

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

ok :)